A positive 20mC charge is placed at the centre of a circle of radius 20 cm. If we move a positive 2mC charge once along the circumference of the circle, will any work be done in the process? Justify your answer.
1
Expert's answer
2018-03-30T09:13:08-0400
Motion of the positive 2 mC charge is going along the equipotential path. Thus, the work done equals to q*(\phi_1-\phi_2) = q*0=0 J.